OpenStudy (anonymous):

Solve for s s -8 = ---- - 10 4

zepdrix (zepdrix):

\[\large\rm -8=\frac{s}{4}-10\]Hey :)

OpenStudy (anonymous):

Hi.

zepdrix (zepdrix):

Remember your PEMDAS thingy? It tells us what order to apply our operations in. Parenthesis taking the highest priority, so it would come first. Multiplication and division before addition subtraction. When we're trying to solve for a variable like s, we would like to do these steps in reverse if possible. That usually ends up being the simplest approach. So in this problem, we have division and subtraction. Let's start by undoing the subtraction. We'll add 10 to each side.\[\large\rm -8+10=\frac{s}{4}-10+10\]giving us,\[\large\rm -8+10=\frac{s}{4}\]What do you get on the left side? c:

OpenStudy (anonymous):

2

zepdrix (zepdrix):

\[\large\rm 2=\frac{s}{4}\]Mmm k good. Now don't think of this as a fraction. Think of the bar as division. How do we undo division? :o what's the opposite of that?

OpenStudy (anonymous):

Multiplication. So s/4 cancels and becomes just s and 2 times 4 is 8 so s=8.

zepdrix (zepdrix):

Yay good job! ୧ʕ•̀ᴥ•́ʔ୨

OpenStudy (anonymous):

LOL. :)

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

Thank you for your help. @zepdrix